Given



1. E = 0.03, 0 =



o = 0.57,



90% confidence



2. E = 0.15, o = 1.2,



95% confidence



3. E = 0.05, o = 0.45,



99% confidence



n=



n=



n=



Solution and Answer

1
Expert's answer
2022-06-14T00:02:12-0400

1.

A 90% confidence interval has a z-score (a critical value) "z_c=1.6449."

"E=z_c\times\dfrac{\sigma}{\sqrt{n}}""n\ge(\dfrac{z_c\sigma}{E})^2"

"n\ge(\dfrac{1.6449(0.57)}{0.03})^2"


"n=977"



2.

A 95% confidence interval has a z-score (a critical value) "z_c=1.96."

"E=z_c\times\dfrac{\sigma}{\sqrt{n}}""n\ge(\dfrac{z_c\sigma}{E})^2"

"n\ge(\dfrac{1.96(1.2)}{0.15})^2"




"n=246"


3.

A 99% confidence interval has a z-score (a critical value) "z_c=2.5758."

"E=z_c\times\dfrac{\sigma}{\sqrt{n}}""n\ge(\dfrac{z_c\sigma}{E})^2"

"n\ge(\dfrac{2.5758(0.45)}{0.05})^2"




"n=538"